OpenStudy (anonymous):

Find two numbers with a sum of –30 and a difference of 8.

OpenStudy (tylerd):

X+Y= -30 X-Y = 8

OpenStudy (anonymous):

A) –20 and –12 B) –18 and –10 C) 19 and 11 D) –19 and –11 ??

OpenStudy (tylerd):

try plugging in for example X+Y= -30 or -20+(-12)=--20-12=-32 so we know right there the answer cant be A) -20 and -12

OpenStudy (anonymous):

oooooh okay. It would be D. !!!! (:

OpenStudy (tylerd):

yep

OpenStudy (anonymous):

Thank you!